OpenStudy (anonymous):

derivative of ((t^7+1)/(t^7-1))^(1/9)

OpenStudy (anonymous):

applying power rule get [(1/9)*(original fraction)^(-8/9)]*(quotient rule on original fraction) ~~mehh ugly, yuck.

OpenStudy (anonymous):

\[\frac{14 t^6 \left(\frac{1+t^7}{-1+t^7}\right)^{1/9}}{9-9 t^{14}} \]